A force of 10N is angle 30degree from x-axis. Find out it's magnitude y -axis direction.
1
Expert's answer
2020-10-07T10:01:32-0400

"\tan \theta ={x \over y}"

where "\theta" is angle between the force and the x-axis

"\tan 30=10/y"

y=17.321N

direction is given by;

"\tan ^{-1} =y/x"

"\theta =\tan ^{-1}\dfrac{17.321}{10}"

=60.00070462

="60\degree"